# Learning Latent Transmission and Glare Maps for Lens Veiling Glare Removal

### v1

@@ 3.5 Stage III: Reversible Restoration
To tackle the spatial-varying veiling glare that hinders blind restoration, we develop DeVeiler, a reversible restoration network that predicts latent maps to invert the underlying degradation. The central component is the Veiling Glare Encoder (VG-Enc), which predicts latent maps $\hat{c}_{vg}=\text{VG-Enc}(I_{de})$ representing the spatial distribution of glare-related parameters. Simply using the predicted maps $\hat{c}_{vg}$ in a one direction manner is ineffective due to a fundamental domain mismatch. Latent maps optimized for conditioning iterative diffusion sampling do not transfer well to guiding the restoration process (see in Tab. 5). To bridge this gap, we introduce a reversibility-guided training paradigm that provides indirect but powerful supervision by leveraging the frozen DDN as a forward degradation model to enforce cycle consistency (Fig. 3(b)). This constraint drives $\hat{c}_{vg}$ to acquire meaningful physical interpretability, as it must parameterize two mutually inverse operations. Within the frozen DDN, the VGIM module uses $\hat{c}_{vg}$ to impose forward modulation, applying the glare degradation to clean features. Symmetrically, in DeVeiler, the VGCM applies an inverse modulation using the same latent maps $\hat{c}_{vg}$ to remove the glare components from degraded features. This structural coupling of VGIM and VGCM encourages the restoration to be a consistent inverse of the learned forward degradation, rather than a loosely correlated statistical mapping.
@@
Fast two-step[ECCV22] [eboli2022fast] 16.62 (73.5%) 0.539 (35.3%) 0.674 (61.3%) 14.94 (78.3%) 0.610 (22.3%) 0.504 (40.3%) 0.390
SwinIR[ICCVW21] [liang2021swinir] 18.18 (62.0%) 0.686 (6.27%) 0.298 (12.4%) 19.34 (40.2%) 0.722 (3.32%) 0.354 (15.0%) 0.782
NAFNET[ECCV22] [chen2022simple] 18.75 (56.7%) 0.684 (6.58%) 0.363 (28.1%) 18.91 (45.8%) 0.723 (3.18%) 0.377 (20.2%) 0.353
DiffBIR[ECCV24] [lin2024diffbir] 17.95 (63.9%) 0.621 (17.4%) 0.398 (34.4%) 18.70 (48.4%) 0.625 (19.4%) 0.412 (26.9%) 66.13
@@
SwinIR[ICCVW21] [liang2021swinir]+RIDCP[CVPR23] [wu2023ridcp] 20.06 (41.4%) 0.695 (4.89%) 0.385 (32.2%) 18.66 (48.8%) 0.692 (7.80%) 0.456 (34.0%) 1.240
SwinIR[ICCVW21] [liang2021swinir]+DiffDehaze[CVPR25] [wang2025learning] 19.31 (50.7%) 0.642 (13.6%) 0.347 (24.8%) 18.28 (53.1%) 0.603 (23.7%) 0.392 (23.2%) 92.54
SwinIR[ICCVW21] [liang2021swinir]+Flare7K++[TPAMI2024] [dai2024flare7k++] 21.67 (15.1%) 0.723 (0.86%) 0.297 (12.1%) 20.74 (17.4%) 0.745 (0.13%) 0.336 (10.4%) 0.990
@@
CycleGAN[ICCV17] [CycleGAN2017] 18.20 (61.8%) 0.558 (30.6%) 0.649 (59.8%) 17.80 (58.0%) 0.623 (19.7%) 0.485 (37.9%) 0.208
UCL-Dehaze[TIP24] [wang2024ucl] 18.55 (58.6%) 0.581 (25.5%) 0.554 (52.9%) 18.69 (48.5%) 0.688 (8.43%) 0.451 (33.3%) 0.208
DiffDehaze[CVPR25] [wang2025learning] 18.87 (55.4%) 0.657 (11.0%) 0.435 (40.0%) 18.55 (50.1%) 0.645 (15.7%) 0.426 (29.3%) 91.76
QDMR[OLT2025] [jiang2025qdmr] 18.45 (59.5%) 0.681 (7.05%) 0.291 (10.3%) 20.67 (18.7%) 0.725 (2.90%) 0.315 (4.44%) 0.428
@@ 4.1 Experiment Settings
Datasets. We verify our method on two distinct, independently evaluated optical systems: a large-aperture single lens (SL) and a metasurface-refractive hybrid lens (MRL). We first constructed a source domain of paired, aberration-only data. Lacking the precise optical parameters required for simulation, we follow prior work [chen2024hybrid, zhang2025degradation] and adopt a screen-capture setup to re-capture clean images from the DIV2K dataset [timofte2017ntire], yielding $170$ (SL) and $125$ (MRL) pairs. We then collected two distinct Target Domains for adaptation, adhering to a strict “Source + one Target” setup. The Screen-Compound domain used the same screen-capture method but with artificial lighting to induce compound degradation, providing GT pairs for $50$ training and 42/2542{/}25 (SL/MRL) test images. The Realworld-Compound domain consists of real-world captures exhibiting natural compound degradation with no GT, providing $50$ training and 51/1151{/}11 (SL/MRL) test images. Finally, we use $500$ clean images from the Flickr2K dataset [timofte2017ntire] to generate $500$ corresponding target-domain degraded pairs. All data across both systems are captured at a resolution of $1280{\times}1920$.
@@
Fast two-step [eboli2022fast] 0.356/2.942/5.821 0.388/2.687/5.352
SwinIR [liang2021swinir] 0.424/3.518/5.710 0.374/3.191/6.696
NAFNET [chen2022simple] 0.374/3.598/5.535 0.398/3.238/5.512
DiffBIR [lin2024diffbir] 0.404/3.476/5.770 0.367/2.761/6.251
@@
SwinIR [liang2021swinir]+RIDCP [wu2023ridcp] 0.386/3.755/5.810 0.399/3.290/6.502
SwinIR [liang2021swinir]+DiffDehaze [wang2025learning] 0.573/3.679/6.141 0.437/3.542/5.230
SwinIR [liang2021swinir]+Flare7K++ [dai2024flare7k++] 0.459/3.528/5.622 0.387/3.319/6.336
@@
CycleGAN [CycleGAN2017] 0.451/2.961/5.300 0.418/3.319/3.836
UCL-Dehaze [wang2024ucl] 0.297/2.763/8.322 0.343/2.963/5.644
DiffDehaze [wang2025learning] 0.406/3.982/6.476 0.428/3.476/6.802
QDMR [jiang2025qdmr] 0.405/3.864/4.773 0.376/3.337/5.509
@@
Implementation Details. VeilGen builds upon pre-trained Stable Diffusion v2-1 [rombach2022high] and is optimized with AdamW [Ilya2019adamw] at a learning rate of $1{\times}10^{-5}$ for 9k iterations with a batch size of 16. The probability parameter $p$ is set to $0.3$ and the mixture coefficient $w$ to $0.85$, consistent with common practices [wang2025learning]. $t^{*}$ is set to 0. DeVeiler is trained in two phases using Adam [kingma2014adam]: Phase I pre-trains on source-domain pairs, and Phase II fine-tunes on a hybrid dataset of $500$ synthetic pairs and source pairs. Each encoder and decoder consists of three groups of ResBlocks, while the bottleneck is constructed with RSTB layers from SwinIR [liang2021swinir]. The balance weight $\lambda_{\text{rev}}$ is set to $1.0$. All experiments are conducted on an NVIDIA A100 GPU. More details are delivered in the supplementary material.
@@ 4.2 Comparison with State-of-the-Art Methods
We compare against three pipelines in Tab. 1 and Tab. 2. First, we evaluate single-degradation baselines, which include the blind CAC method Fast two-step [eboli2022fast] as well as well-known image restoration models (SwinIR [liang2021swinir], NAFNet [chen2022simple], DiffBIR [lin2024diffbir]) re-trained on source domain pairs. Second, cascaded pipelines combine the re-trained SwinIR with pre-trained dehazing (RIDCP [wu2023ridcp], DiffDehaze [wang2025learning]) or flare removal (Flare7K++ [dai2024flare7k++]) models to validate the distinct nature of veiling glare. Finally, the Domain Adaptation (DA) methods are all retrained on our source and target domains for a fair comparison. This includes CycleGAN [CycleGAN2017], UCL-Dehaze [wang2024ucl], the aberration-specific method QDMR [jiang2025qdmr], and a re-trained DiffDehaze [wang2025learning].
Results on Screen-Compound Domain. We first conduct a quantitative evaluation on the Screen-Compound domain, which provides paired GT. We employ full-reference metrics: PSNR and SSIM [wang2004image] to measure fidelity, and LPIPS [zhang2018unreasonable] to assess perceptual quality. As shown in Tab. 1, our DeVeiler significantly outperforms all baselines across all metrics. Single degradation pipeline, trained only on the source domain, fails to generalize to the unseen veiling glare. Cascaded pipelines also prove insufficient. We use the strongest single-degradation model, SwinIR, as the CAC base, followed by pre-trained dehazing or flare removal models. These naive cascades yield unsatisfactory results: dehazing models severely harm color fidelity, while the flare removal model shows negligible perceptual gains. This confirms the distinct nature of veiling glare compared to these other degradations. Furthermore, the DA pipeline produces sub-optimal results due to its physically-agnostic designs. In contrast, DeVeiler leverages the latent veiling glare maps ($c_{vg}$) via the mutually inverse VGIM/VGCM structure, enabling it to precisely reverse the degradation.
@@
Results on Realworld-Compound Domain. We further evaluate the performance of DeVeiler on the challenging Realworld-Compound domain using both SL and MRL systems. As no GT is available, we employ three widely-used no-reference IQA metrics: CLIPIQA [wang2023exploring], Q-Align [wu2023q], and NIQE [mittal2012making]. Tab. 2 shows that DeVeiler achieves excellent performance across the majority of metrics, verifying its generalizability in handling complex, real-world degradations. In addition, we provide visual comparisons in Fig. 5. DeVeiler consistently restores fine details and color fidelity while introducing minimal artifacts, whereas other methods struggle with residual glare and artifacts. Notably, the first row example highlights that aberration and veiling glare are a more pervasive challenge in compact systems than severe flare or ghosting, which are situational. These experimental results demonstrate the robustness of DeVeiler in realistic scenarios. We refer the reader to the supplementary material for more visual results.
@@ 4.3 Ablation Studies
Effectiveness of VeilGen. We compare VeilGen with competitive data generation methods (CycleGAN [CycleGAN2017], DegradationTransfer [chen2021extreme_quality], HazeGen [wang2025learning]). For a fair comparison, an identical restoration network (DeVeiler (w/o VGCM)) is trained on the data generated by each method. As illustrated in Tab. 4, training the restoration network on data from VeilGen yields superior performance, indicating that its generated pairs more accurately capture the target degradation characteristics. The visual comparison in Fig. 6 further shows that VeilGen synthesizes realistic compound degradations while other methods struggle to reproduce the complex degradations.
@@
Analysis of the VGIM/VGCM Mechanism. Fig. 8(b)-(f) visualizes internal feature maps of DeVeiler. The network accurately estimates latent components, identifying low-transmittance and high-glare regions in the input. After modulation by the VGCM, activations in the glare-affected regions are effectively suppressed. This verifies that DeVeiler learns an interpretable restoration process, where the VGCM module specifically targets and removes the glare component, rather than performing a black-box mapping.

### v2

@@ 3.5 Stage III: Reversible Restoration
To tackle the spatial-varying veiling glare that hinders blind restoration, we develop DeVeiler, a reversible restoration network that predicts latent maps to invert the underlying degradation. The central component is the Veiling Glare Encoder (VG-Enc), which predicts latent maps $\hat{c}_{vg}=\text{VG-Enc}(I_{de})$ representing the spatial distribution of glare-related parameters. Simply using the predicted maps $\hat{c}_{vg}$ in a one-direction manner is ineffective due to a fundamental domain mismatch. Latent maps optimized for conditioning iterative diffusion sampling do not transfer well to guiding the restoration process (see in Tab. 5). To bridge this gap, we introduce a reversibility-guided training paradigm that provides indirect but powerful supervision by leveraging the frozen DDN as a forward degradation model to enforce cycle consistency (Fig. 3(b)). This constraint drives $\hat{c}_{vg}$ to acquire meaningful physical interpretability, as it must parameterize two mutually inverse operations. Within the frozen DDN, the VGIM module uses $\hat{c}_{vg}$ to impose forward modulation, applying the glare degradation to clean features. Symmetrically, in DeVeiler, the VGCM applies an inverse modulation using the same latent maps $\hat{c}_{vg}$ to remove the glare components from degraded features. This structural coupling of VGIM and VGCM encourages the restoration to be a consistent inverse of the learned forward degradation, rather than a loosely correlated statistical mapping.
@@
Fast two-step[ECCV22] [11] 16.62 (73.5%) 0.539 (35.3%) 0.674 (61.3%) 14.94 (78.3%) 0.610 (22.3%) 0.504 (40.3%) 0.390
SwinIR[ICCVW21] [29] 18.18 (62.0%) 0.686 (6.27%) 0.298 (12.4%) 19.34 (40.2%) 0.722 (3.32%) 0.354 (15.0%) 0.782
NAFNET[ECCV22] [1] 18.75 (56.7%) 0.684 (6.58%) 0.363 (28.1%) 18.91 (45.8%) 0.723 (3.18%) 0.377 (20.2%) 0.353
DiffBIR[ECCV24] [30] 17.95 (63.9%) 0.621 (17.4%) 0.398 (34.4%) 18.70 (48.4%) 0.625 (19.4%) 0.412 (26.9%) 66.13
@@
SwinIR[ICCVW21] [29]+RIDCP[CVPR23] [56] 20.06 (41.4%) 0.695 (4.89%) 0.385 (32.2%) 18.66 (48.8%) 0.692 (7.80%) 0.456 (34.0%) 1.240
SwinIR[ICCVW21] [29]+DiffDehaze[CVPR25] [49] 19.31 (50.7%) 0.642 (13.6%) 0.347 (24.8%) 18.28 (53.1%) 0.603 (23.7%) 0.392 (23.2%) 92.54
SwinIR[ICCVW21] [29]+Flare7K++[TPAMI2024] [10] 21.67 (15.1%) 0.723 (0.86%) 0.297 (12.1%) 20.74 (17.4%) 0.745 (0.13%) 0.336 (10.4%) 0.990
@@
CycleGAN[ICCV17] [66] 18.20 (61.8%) 0.558 (30.6%) 0.649 (59.8%) 17.80 (58.0%) 0.623 (19.7%) 0.485 (37.9%) 0.208
UCL-Dehaze[TIP24] [51] 18.55 (58.6%) 0.581 (25.5%) 0.554 (52.9%) 18.69 (48.5%) 0.688 (8.43%) 0.451 (33.3%) 0.208
DiffDehaze[CVPR25] [49] 18.87 (55.4%) 0.657 (11.0%) 0.435 (40.0%) 18.55 (50.1%) 0.645 (15.7%) 0.426 (29.3%) 91.76
QDMR[OLT2025] [22] 18.45 (59.5%) 0.681 (7.05%) 0.291 (10.3%) 20.67 (18.7%) 0.725 (2.90%) 0.315 (4.44%) 0.428
@@ 4.1 Experiment Settings
Datasets. We verify our method on two distinct optical systems: a large-aperture single lens (SL) and a metasurface-refractive hybrid lens (MRL). We first constructed a source domain of paired, aberration-only data. Lacking the precise optical parameters required for simulation, we follow prior work [3, 62] and adopt a screen-capture setup to re-capture clean images from the DIV2K dataset [44], yielding $170$ (SL) and $125$ (MRL) pairs. We then collected two distinct Target Domains for adaptation, adhering to a strict “Source + one Target” setup. The Screen-Compound domain used the same screen-capture method but with artificial lighting to induce compound degradation, providing GT pairs for $50$ training and 42/2542{/}25 (SL/MRL) test images. The Realworld-Compound domain consists of real-world captures exhibiting natural compound degradation with no GT, providing $50$ training and 51/1151{/}11 (SL/MRL) test images. Finally, we use $500$ clean images from the Flickr2K dataset [44] to generate $500$ corresponding target-domain degraded pairs. All data across both systems are captured at a resolution of $1280{\times}1920$. See the supplement for further data construction details.
@@
Fast two-step [11] 0.356/2.942/5.821 0.388/2.687/5.352
SwinIR [29] 0.424/3.518/5.710 0.374/3.191/6.696
NAFNET [1] 0.374/3.598/5.535 0.398/3.238/5.512
DiffBIR [30] 0.404/3.476/5.770 0.367/2.761/6.251
@@
SwinIR [29]+RIDCP [56] 0.386/3.755/5.810 0.399/3.290/6.502
SwinIR [29]+DiffDehaze [49] 0.573/3.679/6.141 0.437/3.542/5.230
SwinIR [29]+Flare7K++ [10] 0.459/3.528/5.622 0.387/3.319/6.336
@@
CycleGAN [66] 0.451/2.961/5.300 0.418/3.319/3.836
UCL-Dehaze [51] 0.297/2.763/8.322 0.343/2.963/5.644
DiffDehaze [49] 0.406/3.982/6.476 0.428/3.476/6.802
QDMR [22] 0.405/3.864/4.773 0.376/3.337/5.509
@@
Implementation Details. VeilGen builds upon pre-trained Stable Diffusion v2-1 [40] and is optimized with AdamW [32] at a learning rate of $1{\times}10^{-5}$ for 9k iterations with a batch size of 16. The probability parameter $p$ is set to $0.3$ and the mixture coefficient $w$ to $0.85$, consistent with common practices [49]. $t^{*}$ is set to 0. DeVeiler is trained in two phases using Adam [23]: Phase I pre-trains on source-domain pairs, and Phase II fine-tunes on a hybrid dataset of $500$ synthetic pairs and source pairs. Each encoder and decoder consists of three groups of ResBlocks, while the bottleneck is constructed with RSTB layers from SwinIR [29]. The balance weight $\lambda_{\text{rev}}$ is set to $1.0$. All experiments are conducted on an NVIDIA A100 GPU. More details are delivered in the supplementary material.
@@ 4.2 Comparison with State-of-the-Art Methods
We compare against three pipelines in Tab. 1 and Tab. 2. First, we evaluate single-degradation baselines, which include the blind CAC method Fast two-step [11] as well as well-known image restoration models (SwinIR [29], NAFNet [1], DiffBIR [30]) re-trained on source domain pairs. Second, cascaded pipelines combine the re-trained SwinIR with pre-trained dehazing (RIDCP [56], DiffDehaze [49]) or flare removal (Flare7K++ [10]) models to validate the distinct nature of veiling glare. Finally, the Domain Adaptation (DA) methods are all retrained on our source and target domains for a fair comparison. This includes CycleGAN [66], UCL-Dehaze [51], the aberration-specific method QDMR [22], and a re-trained DiffDehaze [49].
Results on Screen-Compound Domain. We first conduct a quantitative evaluation on the Screen-Compound domain, which provides paired GT. We employ full-reference metrics: PSNR and SSIM [52] to measure fidelity, and LPIPS [63] to assess perceptual quality. As shown in Tab. 1, our DeVeiler significantly outperforms all baselines across all metrics. Single degradation pipeline, trained only on the source domain, fails to generalize to the unseen veiling glare. Cascaded pipelines also prove insufficient. We use the strongest single-degradation model, SwinIR, as the CAC base, followed by pre-trained dehazing or flare removal models. These naive cascades yield unsatisfactory results: dehazing models severely harm color fidelity, while the flare removal model shows negligible perceptual gains. This confirms the distinct nature of veiling glare compared to these other degradations. Furthermore, the DA pipeline produces sub-optimal results due to its physically-agnostic designs. In contrast, DeVeiler leverages the latent veiling glare maps ($c_{vg}$) via the mutually inverse VGIM/VGCM structure, enabling it to precisely reverse the degradation.
@@
Results on Realworld-Compound Domain. We further evaluate the performance of DeVeiler on the challenging Realworld-Compound domain using both SL and MRL systems. As no GT is available, we employ three widely-used no-reference IQA metrics: CLIPIQA [48], Q-Align [55], and NIQE [35]. Tab. 2 shows that DeVeiler achieves excellent performance across the majority of metrics, verifying its generalizability in handling complex, real-world degradations. In addition, we provide visual comparisons in Fig. 5. DeVeiler consistently restores fine details and color fidelity while introducing minimal artifacts, whereas other methods struggle with residual glare and artifacts. Notably, the first row example highlights that aberration and veiling glare are a more pervasive challenge in compact systems than severe flare or ghosting, which are situational. These experimental results demonstrate the robustness of DeVeiler in realistic scenarios. We refer the reader to the supplementary material for more visual results.
@@ 4.3 Ablation Studies
Effectiveness of VeilGen. We compare VeilGen with competitive data generation methods (CycleGAN [66], DegradationTransfer [4], HazeGen [49]). For a fair comparison, an identical restoration network (DeVeiler (w/o VGCM)) is trained on the data generated by each method. As illustrated in Tab. 4, training the restoration network on data from VeilGen yields superior performance, indicating that its generated pairs more accurately capture the target degradation characteristics. The visual comparison (Fig. 6) further shows that VeilGen synthesizes realistic compound degradations while other methods fail.
@@
Analysis of the VGIM/VGCM Mechanism. Fig. 8(b)-(e) visualizes internal feature maps of DeVeiler. The network accurately estimates latent components, identifying low-transmittance and high-glare regions in the input. After modulation by the VGCM, activations in the glare-affected regions are effectively suppressed. This verifies that DeVeiler learns an interpretable restoration process, where the VGCM module specifically targets and removes the glare component, rather than performing a black-box mapping.